Background information Anna has been talking with the company’s directors about the future of Frentheway Farm Equipment (FFE). The company has been using outside suppliers for various key components of the company’s farm equipment, including engines. Anna has decided that FFE should consider the purchase of an engine manufacturer to allow FFE to better integrate its supply chain and get more control over engine features. After investigating several possible companies, Anna feels that the purchase of Mantz Engines Inc (MEI), is a possibility. She has asked Aiden Carney to analyze Mantz’s value. Mantz Engines, Inc., is a privately owned company that was founded seven years ago by Sydney and Claire Mantz. The company manufacturers farm engines used primarily in farm equipment. Mantz has experienced rapid growth because of a proprietary technology that increases the fuel efficiency of its engines with very little sacrifice in performance. The company is equally owned by Sydney and Claire Mantz. The original agreement between the siblings gave each 200,000 shares of stock. The Mantz’s have not issued any additional stock nor have they sold any of their stock since founding the company. Anna has asked Aiden to determine a value per share of Mantz stock. Aiden has gathered the following information about engine manufacturing firms that are publicly traded. EPS DPS STOCK PRICE ROE ROA Jensen Motors Corp. $3.50 $2.00 $37.00 12.00% 9.00% Blue Cow Farm Engines Inc. 8.00 4.00 90.00 15.00% 11.00% Horrocks Farm Equipment (0.50) 1.00 30.00 16.00% 13.00% Horrocks Farm Equipment’s negative earnings per share (EPS) was the result of an accounting write-off last year. Without the write-off, EPS for the company would have been $2.25. Last year, Mantz Engines Inc (MEI) had an EPS of $4.00 and paid a dividend to Sydney and Claire Mantz of $300,000 each. The company also had a return on equity of 16%. Anna tells Aiden that a 14% required rate of return should be used to evaluate the value of Mantz. Background information Anna has been talking with the company’s directors about the future of Frentheway Farm Equipment (FFE). The company has been using outside suppliers for various key components of the company’s farm equipment, including engines. Anna has decided that FFE should consider the purchase of an engine manufacturer to allow FFE to better integrate its supply chain and get more control over engine features. After investigating several possible companies, Anna feels that the purchase of Mantz Engines Inc (MEI), is a possibility. She has asked Aiden Carney to analyze Mantz’s value. Mantz Engines, Inc., is a privately owned company that was founded seven years ago by Sydney and Claire Mantz. The company manufacturers farm engines used primarily in farm equipment. Mantz has experienced rapid growth because of a proprietary technology that increases the fuel efficiency of its engines with very little sacrifice in performance. The company is equally owned by Sydney and Claire Mantz. The original agreement between the siblings gave each 200,000 shares of stock. The Mantz’s have not issued any additional stock nor have they sold any of their stock since founding the company. Anna has asked Aiden to determine a value per share of Mantz stock. Aiden has gathered the following information about engine manufacturing firms that are publicly traded. EPS DPS STOCK PRICE ROE ROA Jensen Motors Corp. $3.50 $2.00 $37.00 12.00% 9.00% Blue Cow Farm Engines Inc. 8.00 4.00 90.00 15.00% 11.00% Horrocks Farm Equipment (0.50) 1.00 30.00 16.00% 13.00% Horrocks Farm Equipment’s negative earnings per share (EPS) was the result of an accounting write-off last year. Without the write-off, EPS for the company would have been $2.25. Last year, Mantz Engines Inc (MEI) had an EPS of $4.00 and paid a dividend to Sydney and Claire Mantz of $300,000 each. The company also had a return on equity of 16%. Anna tells Aiden that a 14% required rate of return should be used to evaluate the value of Mantz.

1. Assuming the company continues its current growth rate, what is the value per share of the company’s stock? (Do not round intermediate calculations)

a. What is the company's total earnings?

b. What are the company's payout and retention ratios? Payout Ratio = Retention Ratio = c. Using the retention ratio, calculate the company's growth rate.

1. Assuming the company continues its current growth rate, what is the value per share of the company’s stock? (Do not round intermediate calculations)

a. What is the company's total earnings?

b. What are the company's payout and retention ratios? Payout Ratio = Retention Ratio =

c. Using the retention ratio, calculate the company's growth rate.

d. Now you can value the company using the entire dividend payment. The total value of the company’s equity under these assumptions is:

e. What is the value per share based on your response to part d?

d. Now you can value the company using the entire dividend payment. The total value of the company’s equity under these assumptions is:

e. What is the value per share based on your response to part d?
Answer #1

1a). Total number of shares (n) = 200,000*2 = 400,000

EPS0 = 4 so

total earnings = EPS0*n = 4*400,000 = 1,600,000

1b). Total dividend given = 300,000*2 = 600,000

Payout ratio = dividend/earnings = 600,000/1,600,000 = 37.50%

Retention ratio = 1-payout ratio = 1-37.50% = 62.50%

c). Growth rate (g) = retention ratio*ROE = 62.50%*16% = 10.00%

d). Firm value = Dividend next year(required return - g)

Dividend next year = last year's dividend*(1+g) = 600,000*(1+10%) = 660,000

Firm value = 660,000/(14%-10%) = 16,500,000 or 16.5 million

e). Value per share = total value/n = 16,500,000/400,000 = $41.25 per share